OpenStudy (anonymous):

Q1.... R=3k and rm=3k plz help :'(

OpenStudy (anonymous):

can anyone write exact answer??

OpenStudy (benzo):

q4 Checked answers p1=V*R2/(R1+R2)*(1-e^(-(R1+R2)*t/(R1*R2*C))) p2=(V*R2/(R1+R2))*cos(1/sqrt(L*C)*(t-T)) p3=C*(V*R2/(R1+R2))*sin((t-T)*1/sqrt(L*C))/sqrt(L*C) have fun...

OpenStudy (benzo):

Q5 E V B Y F Z